What is the probability of drawing a ten or a heart?
Let A= drawing a ten
B= drawing a heart
P(A or B) = P(A) + P(B) - P(A & B)
A & B denotes a ten of hearts.
P( ten or a heart)= 4/52 + 13/52 -1/52 =16/52 = 4/13
What is reason why the 2 sides of a coin are called heads and tails?
Well, generally, on most living things, there are 2 sides, the head and the tail, so, naturally, when they put presidents heads on coins, the term "heads and tails" was coined, pardon the pun.
Head's
there is a head on this side of the coin.
tail's
on the majority of the coins their are animals that have tails on them except for the dime; the dime is the newest coin and the person who said we need a 10cent coin was a sailor but he got lost at see 3 days before the government decided to have a 10 cent coin.
The answer is likely to be a little simpler than that. Also, the terminology is used throughout the English-speaking world: not just in the United States.It is 1248/57120 = 13/595 = 0.0218 approx.
A coin is flipped 3 times what is the probability of 2 tails?
Assuming we want two tails exactly, the possible options to get them are: TTH, THT and HTT. They are three choices out of the eight available, which is a probability of 3/8, 0.375 or 37.5%.
The probability of guessing and getting this question right is 1 out of 2. True of False?
True. When there are only two answers, so guessing should give you a 50% or 1 out of 2 chance of getting the right answer.
The chance of picking a heart is 1/4 as there are four suits. The chance of picking an ace is 1/13, as there are 4 aces in a deck of 52. If I want the probability of drawing a heart of an ace, I can add these probabilities together, but I have double counted on possibility, the ace of hearts. So: Probability = 4/52 + 13/52 - 1/52 = 16/52 or 4/13 Probability = 0.308 expressed as a fraction.
The probability of getting at least 1 answer correct = 1 - Probability of getting all answers correct.
So in your case it for be P(at least 1 answer correct) = 1 - 1/256
where 256 is your sample space, |S| = 2^8.
P( at least one card is a heart) = 1 - P( 0 hearts)
P(0 hearts) = (39/52)(38/51)(37/50)(36/49)(35/48)
P( at least 1 heart) = 1- (39/52)(38/51)(37/50)(36/49)(35/48)
=0.778466
there is a 50% chance. That is very obvious
A classic logic fallacy is to assume that a random chance is changed by how often something has already occurred. It sometimes known as the Gambler's Fallacy.
4 in 7 chance
What is the probability of making a Type II error if the null hypothesis is actually true?
zero. We have a sample from which a statistic is calculated and will challenge our held belief or "status quo" or null hypothesis. Now you present a case where the null hypothesis is true, so the only possible error we could make is to reject the null hypothesis- a type I error.
Hypothesis testing generally sets a criteria for the test statistic to reject Ho or fail to reject Ho, so both type 1 and 2 errors are possible.
13 out of 20
How do you find the experimental probability that a particular result will occur?
You carry out the experiment a large number of times. Count the number of times it was carried out (n). Count the number of times in which the particular outcome occurred (x). Then, the experimental probability for that even is x/n.
Suppose you roll one die Find the probability of not rolling a multiple of 2?
one die has 6 numbers or sides, 3 of which are multiples of 2 (2,4,6) and 3 of which are not (1,3,5). so the odds of not rolling a multiple of 2 are: 3 in 6, simplified to 1 in 2, also known as 50%
Since there are 4 choices the probability of guessing any given answer correctly is 1/4 or .25; call this a success and denote it by p The chance of guessing wrong is .75; call this a failure and denote it by q. So the chance of 3 out of 5 correct answers is 5C3xp^3q^(5-3)=10p^3q^2 5C3x(.25)^3(.75)^2 5x4x3/3x2(.15625)(.5625) 10(.12625)(.5625)=.0877891
What is the probability of getting an odd number when rolling a single 6-sided die?
There are 3 odd numbers, 1, 3 and 5 so the probability of getting an odd number is 3 out of 6 or 1/2. This can also be written as .5 of course.
You toss a coin 5 times what is the probability of getting heads 1 time?
You use the Binomial distribution formula
P(1 head out of 5 tosses) = 5C1 (1/2)^1 (1/2)^4 = 5 / 32
The particular student can be in any one of 3 groups and so the probability of being in group 2 is 1/3.
What year was probability discovered?
Probability is a concept that cannot be "discovered." The concept of probability is ingrained in any complex organism's thought process through evolution and has been ever since complex thought evolved. To survive, animals are constantly taking statistics, weighing odds, figuring out what course of action will lead to survival. For example, an owl might hunt at night and decide to target a certain area of hunting ground, deciding that it is more probable that prey might be found in that area.
However, as a formal category of mathematics, theories regarding probabilities can be developed or "discovered." See the related question below.
What is the probability of getting an even number when throwing a die?
Divide the number of possible even numbers by all the possible numbers. So there are three possible even numbers; 2, 4 & 6. And there are six possible numbers; 1, 2, 3, 4, 5, & 6. So the probability of throwing an even number is 3/6 = .5 .
What are some probability games?
Black Jack, Poker, any dice game, Probhex, however it can also be educational than other probability games.
It is not possible to answer the question because the events for which the probabilities are required, are not defined.